OpenStudy (anonymous):

Given a slope,m, is -3, and passes through point (-1,7) the graph will:

undeadknight26 (undeadknight26):

what do u mean by that? That graph will be?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-1\quad ,&7) \end{array} \\\quad \\ slope = m= -3 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\\ \qquad \qquad \textit{then graph that by picking 2 values for "x"}\)